OpenStudy (anonymous):

how many radians are in 1/8 rotation

Parth (parthkohli):

There are \(\pi\) radians in 1/2 rotations. So, in \(\frac{1}{2}\times \frac{1}{4}\) rotations, there are \(\frac{1}{4} \times \pi\) radians.